OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x − 2y = 6. x-intercept: (0, −3), y-intercept: (0, 2) x-intercept: (−3, 0), y-intercept: (2, 0) x-intercept: (0, 2), y-intercept: (0, −3) x-intercept: (2, 0), y-intercept: (0, −3)

OpenStudy (anonymous):

plug in 0 for x when solving for y, plug in 0 for y when solve for x.

OpenStudy (anonymous):

so is y= -3?

OpenStudy (anonymous):

3(0)-2y=6 -2y=6 so yes

OpenStudy (anonymous):

and x is 2

OpenStudy (anonymous):

yes. :)

OpenStudy (anonymous):

thx so much your help is amazing

OpenStudy (anonymous):

no problem :)

OpenStudy (anonymous):

remember that x=2, so (2,0) and y=-3 so it's (0,-3)

OpenStudy (anonymous):

because you're still using the points, but you know the x value and y value.